A sample of monoatomic hydrogen gas contains 100 atoms and all the atoms are in nth excited state As the atoms come down to the ground state following various transitions, they release a total energy of 1500 Rch IRch -13.6eV]. The maximum possible number of photons emitted in the process s 49 A) 150 C) 4200 B) 2100 D) 6300
Answer #1

Total energy emitted in transition from n th excited state to ground state by one atom = 13.6(1 - 1/(n+1)^2 )

Hence 48/49 = ( 1 - 1/(n+1)^2 )

n = 6

Maximum possible photons emitted by an atom, as it comes down to ground state from n th excited state

= n(n+1)/2 = 21

Maximum possible photons emitted by 100 atoms = 2100

Option b is correct.